The length of the chord joining the points (4cosalpha,4sinalpha) and (4cos(alpha+60°)),4sin(alpha+60°)) of the circle x√2+y√2=16 is

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
These two points lie on a circle.
 Make lines joining origin and each of these points .
Let these lines be OA and OB .
Angle b/w OA and OB = 60°
Each OA and OB make an angle 30° with the y-axis .
Let AB intersect Y axis at C
In right triangle OCA , AC = 4sin 30 = 2
Similarly in right triangle OBA, BC = 2
AB = 2+2 =4 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)